Hi i would like some help on solving these two simultaneous equations for my homework as i am struggling to solve them please help me as it would be very appreciated the questions are
1. one side of the rectangle is 2cm longer than the other. The perimeter of the rectangle is 24 cm. How long are the sides of the rectangle?
2. A fitness fantastic swims 40 lengths of his pool per day. After 5 years, he has swum 1825 km. How long is his swimming pool in meters? (assume there are 365 days per year)
I would be most gratefull if you can help me thank you

Answer by uma(370)

1) let ons side of the rectangle be x cm
The other side = x + 2 cm
Perimeter given = 24 cm
==> 2(sum of the sides of the rectangle) = 24
==> 2(x + x + 2) = 24
==> 2(2x + 2) = 24
==> 4x + 4 = 24
==> 4x = 4 - 4 = 24
==> 4x = 20
==> 4x/4 = 20/4
==> x = 5
Hence the sides of the rectanle measure 5 cm and 7 cm

2) Let the length of the pool be x meters.
Distance swam by him everyday = 40 lengthe
= 40x m
Distance swam in a year = 40x * 365
= 14600x m
Distance swam in 5 years = 14600x*5
= 73000x m
==> 73000x = 1825 km
==> 73000x = 1825000 m
==> 73000x/73000 = 1825000/73000
==> x = 25
So length of the swimming pool = 25 meters

Good Luck!!!